### v3

# VisionDirector: Vision-Language Guided Closed-Loop Refinement for Generative Image Synthesis

Abstract
Generative models can now produce photorealistic imagery, yet they still struggle with the long, multi-goal prompts that professional designers issue. To expose this gap and better evaluate models’ performance in real-world, we introduce Long Goal Bench (LGBench), a 2,000-task suite (1,000 T2I, 1,000 I2I) whose average instruction contains 18—22 tightly coupled goals spanning global layout, local object placement, typography, and logo fidelity. We find even state-of-the-art models satisfy fewer than 72% of the goals and routinely miss localized edits, confirming the brittleness of current pipelines. To address this, we present VisionDirector, a training-free, vision-language supervisor that (i) extracts structured goals from long instructions, (ii) dynamically decides between one-shot generation and staged edits, (iii) runs micro-grid sampling plus semantic verification/rollback after every edit, and (iv) logs goal-level rewards. We further fine-tune the planner with Group Relative Policy Optimization, yielding shorter edit trajectories (3.1 vs. 4.2 steps) and stronger alignment. VisionDirector achieves new state of the art on GenEval (+7% overall) and ImgEdit (+0.07 absolute) while producing consistent qualitative improvements on typography, multi-object scenes, and pose editing.
Figure 1: VisionDirector is a framework that utilizes the VLM Planner to decompose tasks into multiple goals, perform planning and judgment, and progressively optimize both image-editing and image-generation tasks. It achieves performance comparable to, and in some cases even surpassing, closed-source commercial models.
22footnotetext: #Project Lead.
1 Introduction
The field of visual content creation has advanced rapidly with the rise of diffusion models, which now enable a wide range of applications in image generation and editing. However, professional designers still struggle to achieve faithful results because real-world design briefs often consist of long paragraphs with multiple goals, combining global art direction with numerous local constraints on characters, lighting, typography, and logo placement. However, existing benchmarks such as DrawBench[27], TIFA[17], and MagicBrush[40] predominantly use simplified prompts with at most one or two explicit goals, thereby masking model brittleness in realistic, multi-objective scenarios.
To expose this gap and better evaluate models’ performance in real-world production settings, we introduce LongGoalBench (LGBench), which consists of 2,000 tasks sourced from artist workflows and includes multiple real-world targets per prompt. Specifically, each text-to-image (T2I) prompt specifies an average of 18.0 goals across 200 categories and 418 subcategories, for example, nature landscapes and historical scene, while the image-to-image (I2I) split, based on real source photos, contains an average of 11.2 edit directives. We evaluate several state-of-the-art models, including Flux-Krea [20], Flux-Kontex [19], and Qwen-Image [31], on our proposed LGBench. As shown in Table 2, even the leading generative models, such as Flux-Kontex and Qwen-Image, only achieves 55.9% and 71.8% finish score, respectively. And they fail even more frequently on local text and logo edits. These results reveal why current models often fall short in real-world production settings. Although they demonstrate strong aesthetic abilities and can effectively handle single-goal instructions, they struggle with long and multi-goal instructions. Hence, a natural question arises: how can we enhance their ability to process multi-goal instructions without compromising their aesthetic quality?
To solve this question, we propose VisionDirector, a training-free, VLM-guided [36, 8] controller that decomposes long prompts into structured goals, decides whether to solve them in one shot or via staged micro-edits, and verifies/rolls back every edit. Instead of reporting a single reward at the end of generation, VisionDirector inspects each iteration, maintains goal-level state, and emits short instructions to the underlying diffusion models. This design lets us keep high-quality pretrained editors intact while adding a reasoning layer on top. Nevertheless, the controller can still be refined: we further fine-tune the planner with Group Relative Policy Optimization (GRPO), letting the agent experience thousands of long-goal rollouts and quickly learn when to STOP, VERIFY, or EDIT. By applying our proposed VisionDirector framework in post-training, the model is able to generate the desired images more quickly, reducing the average number of generation rounds from 4.2 to 3.1, achieving about a 26% reduction.
Our contributions are three-fold:
  • Benchmark. To bridge the gap between traditional benchmarks and real-world design workflows, we propose LongGoalBench, a benchmark comprising 2,000 long-form tasks with 29,252 annotated goals, dual T2I/I2I modalities, and automated goal-level verification. We release all prompts, reference images, and evaluation scripts to facilitate reproducible future research.
  • Framework. VisionDirector provides a modular closed loop consisting of instruction intake, goal planning, one-shot gating, micro-grid execution, VLM verification/rollback, and decision policies. It improves multi-goal adherence without retraining diffusion backbones.
  • RL fine-tuning. We adapt GRPO to this visual editing context by introducing token-level masking, an alignment-based reward, and rollout workers that render images for every sampled policy. The RL-tuned planner reduces editing rounds by roughly 26% while boosting GenEval and ImgEdit scores.
Taken together, LGBench exposes the challenging cases that current models overlook, VisionDirector addresses these challenges through a director-style agent, and GRPO further refines the policy to produce reliable, instruction-following imagery.
2 LGBench: LongGoal Benchmark
2.1 Benchmark Overview
Figure 2: Distribution of goal types in LGBench for both T2I and I2I subsets. The benchmark balances additive, stylistic, and semantic directives, reflecting the multi-constraint nature of real design tasks.
To bridge the gap between the complex, multi-goal requirements of real-world design workflows and the limited scope of existing benchmarks, we introduce LongGoalBench (LGBench). Unlike prior datasets that evaluate single attributes in isolation, LGBench simulates the actual challenges creative teams face when composing hero shots or refining marketing assets, where success depends on meeting dozens of stylistic, structural, and semantic goals simultaneously.
Scale and coverage.
As shown in Figure 2 and Table 1, LGBench contains $2000$ tasks in total—$1000$ for text-to-image (T2I) generation and $1000$ for image-to-image (I2I) editing—covering 29,25229{,}252 individual goal statements. The T2I subset spans $200$ categories and $418$ fine-grained subcategories (e.g., fantasy_creatures, urban_architecture), averaging $18.0$ goals per prompt (range 15152323). Goal types are balanced across additive objects (31.8%), textual overlays (16.8%), visual effects (16.6%), color constraints (11.8%), and lighting directives (11.5%), ensuring diverse coverage of stylistic and compositional aspects. The I2I subset focuses on realistic photo edits across $29$ coarse classes (densely covering animals, people, and nature), mapped to roughly $710$ subcategories. Each I2I instruction includes $11.2$ goals on average (range 10102222), with emphasis on effects (34.4%), color grading (21.5%), typography (17.5%), and lighting refinement (15.8%).
2.2 Benchmark Construction
Figure 3: Construction pipelines for LongGoalBench.
To scale dataset creation beyond hand-written prompts, we employ a semi-automated pipeline powered by Claude 4.5 Sonnet through our VisionDirector framework showing in Figure 3. Claude serves as a prompt composer: it reads structured goal inventories, reasons about dependencies, and produces coherent, machine-verifiable instructions in natural language.
T2I generation pipeline. (1) Goal pool preparation. For each of the 200 high-level categories (e.g., fantasy creatures, urban architecture), we define a concise inventory of fine-grained visual goals such as “add molten-gold pile at high prominence” or “set color temperature to 3200 K”. Each goal is annotated with a numeric strength (0–100%) and a goal type (object, text, color, lighting, or effect).
(2) Prompt composition via Claude. Claude receives the goal list together with category metadata, then organizes and merges all goals into a single coherent instruction that includes quantitative cues (e.g., “at moderate intensity (45%)”) while avoiding contradictions. Each resulting prompt record contains the category labels, the composed long-form instruction, and structured annotations for individual goals.
(3) Statistics. The final 1,000 T2I prompts span 418 subcategories and include $18\pm 3$ goals per prompt, contributing approximately 18k individual objectives.
I2I editing pipeline. (1) Seed image generation. We define 29 coarse scene classes covering people, animals, and nature, and generate a representative base image for each using the Flux Krea model as the editable “before” reference.
(2) Edit instruction authoring. Claude (vision-enabled) is provided with the base image and its class description. It analyzes salient regions—for instance, detecting faces, lighting, or text overlays—and formulates 10–22 concrete edit directives such as “increase hair shine by 60%” or “overlay TIMELESS BEAUTY in elegant serif font”. Because Claude observes the actual image, the edits it proposes remain spatially and semantically consistent with the layout.
(3) Data structure. Each I2I instruction entry stores the class metadata, long-form directive, and structured goal descriptors, while the paired source images are linked via unique identifiers. The I2I subset contributes 11,217 additional goals, yielding a total of 29,252 annotated objectives across both modalities.
Metric T2I I2I
Overall Statistics
Total Samples 1,000 1,000
Total Categories 200 29
Total Subcategories 418 710
Avg. Prompt Length 207.6 253.4
Avg. Goals/Sample 18.04 11.22
Total Goals 18,035 11,217
Goal Type Count
Add Object 5,739 760
Text 3,032 1,964
Effect 2,993 3,863
Color 2,137 2,409
Lighting 2,069 1,773
Composition 2,065 448
Table 1: Statistics result of the LGBench.
[…]
Base models (no VisionDirector)
Qwen-Image (T2I)  [31] T2I 71.8 30.7 76.6 29.7 87.0 78.3 85.8 77.5
Flux-Krea (T2I)  [20] T2I 66.8 18.6 75.7 25.6 82.3 67.7 78.1 67.8
Flux-Dev (T2I)  [19] T2I 40.0 1.0 32.2 0.8 55.4 58.2 62.9 55.1
Qwen-Edit+ (I2I)  [31] I2I 71.0 52.4 62.0 93.2 60.3 77.0 70.9 49.3
Qwen-Edit (I2I)  [31] I2I 65.3 40.1 56.7 66.4 60.1 74.6 69.3 54.1
Flux-Kontext (I2I)  [2] I2I 55.9 31.0 49.7 70.8 48.3 62.6 53.6 39.7
Base + VisionDirector (VDred: absolute gain over base, per goal type)
Qwen-Image + VID T2I 74.8+3.0 36.7+6.0 79.1+2.5 37.2+7.5 88.4+1.4 80.6+2.3 87.3+1.5 79.9+2.4
Flux-Krea + VID T2I 74.9+8.1 35.9+17.3 81.6+5.9 43.7+18.1 86.6+4.3 75.6+7.9 83.4+5.3 75.7+7.9
Flux-Dev + VID T2I 62.4+22.4 25.6+24.6 57.5+25.3 29.5+28.7 75.9+20.5 77.6+19.4 80.1+17.2 74.0+18.9
Qwen-Edit+ + VID I2I 76.7+5.7 63.5+11.1 69.5+7.5 94.5+1.3 68.1+7.8 81.5+4.5 76.6+5.7 59.3+10.0
Qwen-Edit + VID I2I 70.6+5.3 56.4+16.3 63.3+6.6 71.5+5.1 66.2+6.1 78.5+3.9 74.0+4.7 61.1+7.0
Flux-Kontext + VID I2I 62.4+6.5 49.7+18.7 57.1+7.4 75.1+4.3 55.9+7.6 68.1+5.5 60.4+6.8 48.6+8.9
Table 2: Performance comparison of models with and without VisionDirector.
Property LGBench DrawBench [28] TIFA [17] GenEval [13] MagicBrush [40] ImgEdit [37]
Dual Modalities (T2I + I2I)
Long-chain Instructions (10++ goals)
Hybrid-goal Tasks (>5>5 different goals)
Automated Goal-level Verification
Table 3: Comparison between LGBench and other T2I/I2I benchmarks.
Automated execution. All tasks are generated and evaluated automatically through a distributed execution pipeline that assigns workloads across GPUs, loads the designated diffusion or editing backbone (e.g., Flux, Qwen), and enforces consistent resolution and formatting. Each run outputs unified image, metadata, and prompt files in a standardized format. This modular design allows new generative models to be integrated seamlessly without re-annotation or manual intervention.
2.3 Evaluation Results
Each generated image or edited result is automatically evaluated by a multimodal verifier built on Qwen3-VL-32B-Instruct. For image-to-image tasks, the verifier compares the “before” and “after” images; for text-to-image tasks, it inspects the generated output directly. Long-form prompts are internally converted into structured instructions so that the verifier can reason step by step, produce explanations, and return verdicts for each goal. Responses with confidence below 0.81 are filtered out. A task is labeled as success if at least 80% of its goals are satisfied, partial if coverage lies between 0–80%, and failure if no goal passes. We also compute per-goal-type pass rates to diagnose which visual dimensions degrade first.
Across both tracks, Qwen-based generators exhibit the strongest multi-goal consistency. In the text-to-image track, Qwen-Image achieves a 71.8% overall goal success rate, clearing the 80% threshold on nearly one third of all prompts. Flux-Krea trails by roughly five percentage points, while Flux-Dev collapses to 40.0% and almost never satisfies text-related goals (0.8% success). On the editing side, Qwen-Edit+ attains 71.0% goal success with more than half of its tasks fully completed, outperforming both Qwen-Edit (65.3%) and Flux-Kontext (55.9%), the latter showing a 7.3% failure rate.
These trends reveal consistent bottlenecks in typography, additive object placement, and lighting refinement—dimensions often ignored by single-score benchmarks. We plan to extend evaluations to additional state-of-the-art and closed-source systems (e.g., HiDream, Janus, Bagel, Nana Banana, and SeedDream 4.0) once access becomes available. Overall, the results confirm that LGBench effectively exposes where current diffusion and editing models break down under long, interdependent goals, motivating the need for a higher-level director that can reason about multi-step objectives. This insight directly inspires the design of VisionDirector, introduced in the next section.
Figure 4: Workflow of VisionDirector. The planner interprets long instructions, decides between one-shot or staged execution, performs micro-grid sampling, verifies progress, and rolls back if needed.
2.4 Discussion and Motivation
LGBench is designed to evaluate multi-attribute alignment rather than simple prompt fidelity. In real-world creative production, tasks rarely involve the one-shot fulfillment of a single textual cue; instead, they require the coherent realization of multiple interdependent goals such as composition, lighting, color harmony, and textual consistency. However, existing evaluation suites offer limited insight into this regime. As summarized in Table 3, previous benchmarks such as DrawBench [28], and TIFA [17] rely on short prompts and human preference studies. GenEval [13] and MagicBrush [40] introduce a limited number of QA-style or short edit tasks but still emphasize subjective or low-granularity scoring. None provide large-scale, reproducible evaluations in which success depends on meeting 10–23 quantitative goals verified automatically.
LGBench fills this gap by expanding both the instruction space and the scoring granularity. Each task consists of long, structured goal chains authored by Claude, paired with Flux-generated reference images and verified through a Qwen-based vision–language pipeline. This design enables the systematic measurement of fine-grained attributes—for example, assessing whether typography is legible and properly positioned, or whether the light–color balance and overall composition remain consistent after multiple edits. By quantifying such aspects, LGBench reveals nuanced failure modes in modern diffusion and editing systems that were previously anecdotal, such as neglecting embedded text or over-saturating local illumination.
Beyond ranking models, LGBench serves as infrastructure for studying goal-conditioned training, adaptive guidance, and verification-aware generation. Practitioners can examine per-goal statistics to identify where models fail, then refine training data, negative prompts, or guidance schedules accordingly. As shown in Table. 2, even state-of-the-art commercial systems repeatedly miss local text or logo constraints when instructions contain 15 or more interacting goals, underscoring the need for a director-style agent to manage multi-goal consistency. Hence, we try to propose a solution to solve this challenge in Sec. 3.
[…]
Show-o  [34] 0.95 0.52 0.49 0.82 0.11 0.28 0.53
Emu3-Gen [30] 0.98 0.71 0.34 0.81 0.17 0.21 0.54
PixArt-$\alpha$  [6] 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SD3 Medium [10] 0.98 0.74 0.63 0.67 0.34 0.36 0.62
FLUX.1 [Dev]  [19] 0.98 0.81 0.74 0.79 0.22 0.45 0.66
SD3.5 Large  [11] 0.98 0.89 0.73 0.83 0.34 0.47 0.71
JanusFlow  [23] 0.97 0.59 0.45 0.83 0.53 0.42 0.63
Lumina-Image 2.0  [25] 1.00 0.87 0.67 0.93 0.34 0.62 0.73
Janus-Pro-7B  [7] 0.99 0.89 0.59 0.90 0.79 0.66 0.80
HiDream-I1-Full  [4] 1.00 0.98 0.79 0.91 0.60 0.72 0.83
GPT Image 1 [High]  [24] 0.99 0.92 0.85 0.92 0.75 0.61 0.84
Seedream 3.0  [12] 0.99 0.96 0.91 0.93 0.47 0.80 0.84
Qwen-Image  [31] 0.99 0.92 0.89 0.88 0.76 0.77 0.87
VisionDirector 0.99 0.94 0.96 0.92 0.88 0.95 0.94
Table 4: Results on GenEval Bench. “Single Obj.” and “Two Obj.” denote single- and two-object prompts. The “Overall” column averages the scores across the nine tasks.
Figure 5: I2I result comparison with other models.
3 VisionDirector
3.1 Overview
Real-world creative tasks often require iterative coordination rather than one-shot generation. VisionDirector instantiates a closed-loop “director” that plans, executes, and verifies visual edits under long, interdependent instructions. The system integrates three modular components: (1) a planner VLM (Qwen3-VL-8B [1] ) that parses and schedules goals, (2) text-to-image and image-to-image editors (Qwen-Image, Qwen-Image-Edit) that perform visual updates, and (3) a verifier VLM that provides goal-level feedback. The planner owns the semantics: it decomposes the user instruction into structured goals, decides execution order, and supervises revisions until convergence. Editors are stateless executors that follow the planner’s directives, while the verifier evaluates results and signals when to stop. Because each module communicates through natural-language messages, components can be swapped or upgraded without retraining the planner, and the same control policy can later be fine-tuned through reinforcement learning (§3.3).
3.2 Training-Free Closed-Loop Control
Figure 4 illustrates the overall workflow. The training-free controller operates through eight deterministic stages inspired by Deepseek-R1-style [14] reasoning but specialized for visual generation.
Step1: Instruction intake. A raw multimodal instruction (optionally with a reference image) enters the system. Basic metadata such as resolution and language tags are normalized before reasoning.
Step2: Goal extraction and planning. The planner extracts goals verbatim from the instruction, assigning each a type (global, local, text, layout, etc.), a conflict flag, and a scalar score estimating one-shot feasibility. All goals are stored in a pending list; a separate completed set starts empty.
Step3: One-shot or staged execution. If the estimated one-shot score is high and the affected area small, the planner directly attempts a complete composition. Otherwise, it schedules goals in batches of one or two, ordered from global to local constraints, so that each edit remains focused.
Step4: Micro-grid execution. For each active batch, the planner issues a concise directive (e.g., “Raise the key light on the hero’s left cheek”). Multiple candidates are generated with different random seeds, and a lightweight VLM judge selects the best image. This mitigates diffusion randomness without significant latency.
Step5: Verification and rollback. A verifier VLM evaluates results against pending goals, marking each satisfied or not. If the new edit degrades overall alignment, the system reverts to the previous best image and reshuffles remaining goals to avoid repeating the same conflict.
Step6: Decision and continuation. If all goals are satisfied, the loop terminates. Otherwise, the planner adapts the editing mode: local adjustments trigger inpainting; unsatisfied global constraints prompt scene-level regeneration. Every few rounds, a self-query (“Can the image still improve?”) triggers early stopping if confidence is low.
Step7&8: Looping and outputs. The loop continues until no further improvements are detected. The final image and full reasoning trace (goals, edits, verdicts) are returned for transparent evaluation and auditing.
3.3 GRPO for Efficient Generation
While the training-free VisionDirector already enhances instruction following, it still requires multiple rounds of editing before convergence. Furthermore, VisionDirector is not only a training-free pipeline but can also be employed after training to further improve generation efficiency. Specifically, we post-train the planner VLM using Group Relative Policy Optimization (GRPO) [29], a reinforcement learning method that encourages concise, high-reward edit strategies.
Let $x$ denote the multimodal prompt and $y=(y_{1},\ldots,y_{T})$ the planner’s interleaved Describe–Inspect–Revise actions. We train the planner with a GRPO variant of PPO over $G$ sampled trajectories per prompt:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x,\{y^{(i)}\}}\Bigg[$ $\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\sum_{t}I(y^{(i)}_{t})}\sum_{t}I(y^{(i)}_{t})\;\mathcal{L}_{\text{clip}}\!\big(\rho^{(i)}_{t},\,\hat{A}^{(i)}_{t}\big)$ (1)
$\qquad-\;\beta\,\mathrm{KL}\!\big(\pi_{\theta}\,\|\,\pi_{\text{ref}}\big)\Bigg],$
where $I(\cdot)$ masks tokens originating from external tools so that updates apply only to the planner’s textual actions, and $\beta$ controls KL regularization toward a reference policy $\pi_{\text{ref}}$. Here $\mathcal{L}_{\text{clip}}$ is the standard PPO-style clipped surrogate loss, $\rho^{(i)}_{t}$ is the per-token importance ratio, and $\hat{A}^{(i)}_{t}$ is the group-normalized advantage using the average group reward as a baseline.
Reward model and rollout. Each trajectory replays the entire director loop and generates a final edited image. A separate alignment VLM scores how well the image satisfies the target instruction on a 0–5 scale, providing dense pixel-level supervision for attributes such as typography, lighting, and spatial layout. Rewards are normalized within each group before policy updates.
Effect on inference. The GRPO-tuned planner cuts the median editing rounds from 4.2 to 3.1, reducing inference latency by about 26% without lowering verification accuracy. Empirically, the optimized policy proposes higher-gain edits earlier and terminates confidently once only minor residual goals remain, effectively complementing the training-free director pipeline above.
4 Experiments
4.1 Evaluation Setup
Benchmarks. We evaluate VisionDirector on two benchmarks that stress long-horizon instruction following. For text-to-image (T2I), we report accuracy on GenEval [13]. GenEval focuses on compositional prompts with multiple objects and attributes. For image-to-image (I2I) editing we adopt ImgEdit [37], which covers nine edit primitives.
Implementation Details Unless otherwise stated we run the VisionDirector loop with four T2I samples or a single I2I edit per iteration and cap the closed-loop refinement at six iterations. All competing models use their publicly released settings; tables retain the original leaderboard numbers so that improvements attributable to VisionDirector can be read in-context. Appendix material provides additional ablations on iteration count and verifier confidence thresholds.
4.2 Results on Text-to-Image
Table 4 summarizes compositional fidelity on GenEval. VisionDirector, instantiated with a Qwen-Image backbone and the GRPO-trained planning agent, delivers the highest overall score (0.94) while jointly improving all sub-metrics, especially attribute binding and reasoning about relative positions.
4.3 Results on Image-to-Image
For editing tasks we follow ImgEdit. ImgEdit scores nine edit primitives from 1–5 via GPT-4.1 judges. VisionDirector surpasses prior open agents on every ImgEdit primitive (Table 5) indicating that the closed-loop controller faithfully preserves unedited regions while applying instruction-specific changes.
[…]
MagicBrush  [40] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.22 1.90
Instruct-Pix2Pix  [3] 2.45 1.83 1.44 2.01 1.50 1.44 3.55 1.20 1.46 1.88
AnyEdit  [18] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.65 2.45
UltraEdit  [42] 3.44 2.81 2.13 2.96 1.45 2.83 3.76 1.91 2.98 2.70
OmniGen  [33] 3.47 3.04 1.71 2.94 2.43 3.21 4.19 2.24 3.38 2.96
ICEdit  [41] 3.58 3.39 1.73 3.15 2.93 3.08 3.84 2.04 3.68 3.05
Step1X-Edit  [22] 3.88 3.14 1.76 3.40 2.41 3.16 4.63 2.64 2.52 3.06
BAGEL  [9] 3.56 3.31 1.70 3.30 2.62 3.24 4.49 2.38 4.17 3.20
UniWorld-V1  [21] 3.82 3.64 2.27 3.47 3.24 2.99 4.21 2.96 2.74 3.26
OmniGen2  [32] 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 4.68 3.44
FLUX.1 Kontext [Pro]  [2] 4.25 4.15 2.35 4.56 3.57 4.26 4.57 3.68 4.63 4.00
GPT Image 1 [High]  [24] 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.89 4.20
Qwen-Image-Edit  [31] 4.38 4.16 3.43 4.66 4.14 4.38 4.81 3.82 4.69 4.27
VisionDirector 4.55 4.32 3.72 4.83 4.41 4.52 4.92 4.05 4.72 4.35
Table 5: Results on ImgEdit Bench. The scores range from 1 to 5, representing quality from low to high. The “Overall” column averages the scores across the nine tasks. Our VisionDirector significantly outperforms the open-source model and demonstrates competitive performance with the closed-source model.
4.4 Ablations and Analysis
We ablate core design choices on a held-out LGBench validation split of 200 mixed T2I/I2I tasks. Table 6 shows that GRPO shortens trajectories by 26% while improving goal coverage and reducing edit cost. Removing the micro-grid sampler increases stochastic failures on typography instructions by 7%, while disabling semantic rollback leads to accumulated hallucinations and a drop in goal coverage to 0.74 (see supplementary for details).
Figure 6 analyzes VLM decision-making under increasing task complexity. The system operates in a high-efficiency regime for tasks with $\leq$15 goals, requiring only 1–3 iterations. Beyond this point, the preference for one-shot execution sharply declines—from over 85% to below 10% at 30 goals—indicating an adaptive shift toward staged execution for complex multi-goal tasks.
Table 7 further shows that the refinement strategy outperforms alternative methods, and combining strategies yields the best overall performance.
Figure 6: Adaptive decision-making in VisionDirector. The VLM exhibits distinct behavioral phases based on task complexity. (a) Iteration steps increase from 1-3 to 5-7 after exceeding 15 goals. (b) One-shot execution preference (>>85%) transitions to staged execution (<<10%) beyond the critical threshold, demonstrating the system’s adaptive control strategy.
[…]
VisionDirector (no RL) 4.2 0.74 3.3
VisionDirector + GRPO 3.1 0.78 2.5
Table 6: GRPO improves planning efficiency on LGBench. “Goal cov.” measures the fraction of verified goals per task; “Edits” counts diffusion executions.
4.5 Case Study
Qualitative examples for both modalities are highlighted in Figure. 5; higher-resolution versions plus interactive galleries will appear on the project page, and additional breakdowns are deferred to the supplementary PDF.
5 Discussion
5.1 Why we need agent framework for generation?
The results on LGBench demonstrate that the shortcomings of current leading diffusion models do not stem from a lack of texture fidelity or aesthetic quality, but rather from their misinterpretation of long prompts and inability to comprehend complex semantic information.
Our proposed VisionDirector explicitly delegates semantic processing to a planner VLM [35], which extracts goals, reasons about conflicts, and verifies each edit, while the diffusion backbones focus solely on rendering. Through this separation of concerns, the model preserves visual quality, generates detailed logs for debugging, and maintains a modular architecture. Moreover, a more capable editor or verifier can be substituted without retraining the planner.
Method Goal $\geq 80$%
Baseline(Flux-Krea) 66.8 18.6
+ Reprompting 69.0 22.7
+ Best-of-N(N=4) 70.5 23.1
+ Refinement 71.2 29.5
All Strategies 74.2 35.2
Table 7: Ablation study on different optimization strategies (values in %). “Goal” reports overall goal success rate, while $\geq$80% indicates task-level high quality coverage. Each strategy is tested independently except the final row which combines all.
5.2 Broader Impact
Generate Diverse Training Dataset. Due to its high accuracy and fully automated nature, VisionDirector can be used to construct large-scale datasets, which can then be employed to train domain-specific or more advanced generative models. Furthermore, the post-trained VisionDirector model can better adapt to the text distribution requirements of generation models. Compared with manually adjusting images through multiple prompt iterations, it can produce the desired images in fewer rounds without human intervention, significantly reducing token consumption and saving substantial time.
6 Related Work
Diffusion generation and editing. Large diffusion models such as DDPM/Score models [15] and latent diffusion variants [26] have driven recent leaps in visual quality. Follow-up work tailors them for instruction-based editing, e.g., Instruct-Pix2Pix [3] or the more recent unified pipelines (Step1X-Edit, BAGEL, UniWorld) summarized in Table 5. Beyond images, diffusion-based video generation [5] and editing [39, 38] further magnify these limitations, as temporal consistency and long-horizon instruction adherence are difficult to maintain with single-shot prompting. These systems excel at local photorealism but still rely on single-shot prompts; none reason explicitly about dozens of heterogeneous goals. VisionDirector keeps these high-quality editors intact and adds a director layer that decomposes and monitors long instructions.
Benchmarking long-form instructions. Multi-goal evaluation has been explored in compositional prompts (GenEval [13]) and dense textual QA (DPG [16]), but these suites cover at most two goals per prompt and focus on T2I only. ImgEdit [37] and GEdit [22] evaluate instruction-based editing, yet their directives remain short and lack goal-level metadata. LGBench complements these efforts by providing dual modalities, 29k annotated goals, and a verifier that outputs structured pass/fail tags, enabling agent-style analysis.
7 Conclusion
We presented LGBench, a dual-modality benchmark that stresses instruction following with 29k annotated goals, and VisionDirector, a director-style framework that uses VLMs to plan, verify, and refine diffusion outputs. The training-free controller already improves multi-goal fidelity by decomposing prompts, running micro-grid edits, and rolling back regressions; a subsequent GRPO phase further shortens trajectories and boosts quantitative scores across GenEval and ImgEdit. Our study suggests that pairing strong diffusion priors with lightweight reasoning agents is a practical path toward professional-grade creative tooling. Future work will scale the benchmark to video and 3D assets, tighten verifier alignment with human preference data, and explore collaborative editing loops that keep humans in the loop for subjective art direction.